# Modeling dissipation in quantum active matter

### v1

@@ Abstract
Active matter denotes a system of particles immersed in an external environment, from which the particles extract energy continuously in order to perform motion. Extending the paradigm of active matter to a quantum framework requires an open quantum system description. In this work, we consider a driven quantum particle whose external driving exhibits characteristics of classical activity. We model the dynamics with time-local master equations and analyze the particle motion at different time scales for different forms of the master equations. By systematically comparing several types of master equations, we uncover how the particle motion evolves under the interplay of quantum effects and active-like dynamics. These results are essential for guiding possible experiments aimed at realizing quantum analogues of classical active systems.
@@ I Introduction
Quantum active matter is an emerging and rapidly developing field [1, 2, 3, 4, 5, 6, 7] that aims to find a quantum-mechanical particle which can be classified as an active one [8]. In classical active matter, activity arises from the local absorption of energy from internal or external sources and its subsequent conversion into persistent, directed motion [9, 10, 11]. This mechanism intrinsically drives the system far from equilibrium. This definition can be straightforwardly extended to quantum systems. However, it encounters the challenge to identify the conditions that meet the requirements.
@@
In classical active matter, dissipation is a crucial factor since active matter dynamics and thermodynamics typically arise from the energy exchange with an external environment [12, 13, 14, 15]. A quantum mechanical analogue of active matter shall also possess these features and necessarily calls for an adequate formalism, which needs to be developed and verified.
@@
In quantum mechanics, the theory of open quantum systems [16] provides a sound description of the dynamics of a quantum system that exchanges energy and establishes correlations with an external reservoir. In the presence of an external reservoir, the unitary evolution of the state vector observed in isolated (non-dissipative) quantum systems is described by a completely positive and trace-preserving (CPTP) time evolution of the system’s density matrix. The latter determines the evolution of the system’s state and consistently encompasses the quantum probability amplitudes of the quantum dynamics, from which interferences originate, and the classical probability due to partial knowledge about the system, which dominates the dynamics in the classical limit.
@@
Generators of the CPTP map can be cast in the form of Nakajima-Zwanzig master equations [17, 18, 19], which can be rigorously derived by eliminating the variables of an external environment from the equations of motion of the system’s degrees of freedom and are generally nonlinear and inhomogeneous differential equations for density operators. Their specific form depends on the initial state of the reservoir, on the initial correlations between system and reservoir, and on the system-reservoir interactions. These equations are tractable in few limits. One prominent case is found when the Born-Markov approximation holds, for which the master equation becomes a homogeneous differential equation that is local in time. Different types of time-local master equation have been discussed and used. Some master equations are designed to satisfy the CPTP requirement [20, 21, 22]. For instance, secular [23, 16, 24] and rotating-wave [25, 24] approximations are well suited for the description of quantum optical systems. Other approaches are designed to be thermodynamically consistent, making them more natural from the perspective of statistical mechanics [26, 27, 28]. These reproduce the classical Fokker-Planck equation in the classical limit [29, 30]. This distinction is not merely theoretical. Different master equations best describe different experimental scenarios [31, 32]. Understanding how quantum active dynamics can emerge for different dissipators, describing dissipation and dephasing, can be crucial for experiments aiming at observing quantum active dynamics.
@@
The purpose of this work is to analyse the effect of different types of quantum time-local master equations on the dynamics of quantum active matter. Our starting point is the framework of Ref. [5], where the features of active matter reproduced by a quantum particle confined by a potential and driven with colored noise. In this paper we extend this approach to consider different generators of the open-quantum system dynamics. By systematically comparing several dissipative frameworks, we uncover how quantum effects and active-like dynamics interplay differently across weak and strong dissipation regimes. Depending on the underlying dissipator, one may obtain a quantum-active evolution that ensures positivity of the density matrix, or one that reproduces classical active matter dynamics in the classical limit – but not both simultaneously.
II Nonequilibrium dynamics in open quantum system
The purpose of this section is to introduce a basic model of a quantum active particle in the presence of an external environment. We use plausible models of the open quantum system dynamics and analyse when these models do reproduce the classical limit of Ref. [5], where particle’s self-propulsion arises from colored (Ornstein-Uhlenbeck) noise.
@@ II.1 Model
The minimal model of quantum active matter of Ref. [5] is inspired by the active Ornstein-Uhlenbeck particle (AOUP). In the classical AOUP model, the particle’s self-propulsion arises from colored noise $x_{c}(t)$, which drives persistent active motion [33, 34, 35, 36, 37, 38, 39]. Specifically, the variable $x_{c}(t)$ is the trajectory of a classical particle that solves the equation of motion
@@
where $\eta(t)$ is a Gaussian white noise with zero mean and correlations $\overline{\eta(t^{\prime})\eta(t^{\prime\prime})}=\delta(t^{\prime}-t^{\prime\prime})$. Since $x_{c}(t)$ represents the stochastic trajectory of an externally driven classical degree of freedom, physical observables must be averaged not only over quantum states but also over the ensemble of possible realizations of $x_{c}(t)$. In other words, the quantum system is subject to a noisy classical driving, and its effective behavior emerges only after performing both the quantum and the classical ensemble average.
In the quantum-mechanical model proposed in Ref. [5] a quantum particle of mass $m$ is confined by a time-dependent harmonic potential. The potential has fixed frequency $\omega$ but the center follows the prescribed Ornstein-Uhlenbeck trajectory $x_{c}(t)$. Thereby, the quantum particle is governed by a time-dependent Hamiltonian of the form
@@
where $x_{c}(t)$ is a classical variable undergoing the dynamics of Eq. (1); $\hat{x}$ and $\hat{p}$ are the canonically-conjugated position and momentum along $x$, satisfying the commutation relations $\hat{x},\hat{p}]=i\hbar$. In this setting, the motion is one dimensional, and the classical particle establishes a contact between the quantum particle and a thermal bath. Since the external drive $x_{c}(t)$ is governed by by the classical equation of motion (1) and unaffected by the state of the quantum particle, we can represent the state of the quantum system by a probability distribution over quantum trajectories $|\psi(x_{c}(t),t)\rangle$. Here, $|\psi(x_{c}(t),t)\rangle$ is the solution of the Schrödinger equation $\partial_{t}|\psi(x_{c}(t),t)\rangle=\hat{H}(x_{c}(t))|\psi(x_{c}(t),t)\rangle$ for a given trajectory $x_{c}(t)$. For this reason, we can represent the state of the quantum particle by a density matrix $\hat{\varrho}(t)$,
$\hat{\varrho}(t)=\int Dx_{c}\mathcal{P}[x_{c}]|\psi(x_{c}(t),t)\rangle\langle\psi(x_{c}(t),t)|\,$, (3)
where $P[x_{c}]$ denotes the probability functional that the particle follows the trajectory $x_{c}(t)$ and $\int Dx_{c}$ is therefore an integration over all paths, such that $\int Dx_{c}\mathcal{P}[x_{c}]=1$.
In this work we assume that the quantum active particle is additionally coupled to a reservoir. The resulting dynamics results from the interplay of the energy exchange of the quantum particle with an external reservoir and the coupling with the classical active particle, undergoing the dynamics of Eq. (1). For a given trajectory $x_{c}(t)$, the state of the system is now described by a density operator $\hat{\rho}(x_{c}(t),t)$. In general, the state $\hat{\rho}(x_{c}(t),t)$ is connected to the initial state $\hat{\rho}(x_{c}(0),0)$ by the CPTP map $\hat{\rho}(x_{c}(t),t)=\hat{\Lambda}_{x_{c}(t)}[\hat{\rho}(x_{c}(0),0)]$. By averaging over all trajectories, the state of the system is now given by
@@
In order to determine the dissipative dynamics, we consider generators that are local in time, such that $\partial_{t}\hat{\rho}=\hat{\mathcal{L}}(x_{c}(t),t)[\hat{\rho}$]. Here, $\hat{\mathcal{L}}$ has the form
$\hat{\mathcal{L}}(x_{c}(t),t)[\hat{\rho}$]
$=$ $-\frac{i}{\hbar}\left[\hat{H}(x_{c}(t)),\hat{\rho}(x_{c}(t),t)\right]+\hat{\mathcal{D}}(x_{c}(t),\hat{\rho}(x_{c}(t),t)),$
where $\left[\cdot,\cdot\right]$ is the commutator $\left[\hat{A},\hat{B}\right]=\hat{A}\hat{B}-\hat{B}\hat{A}$ and $\hat{\mathcal{D}}$ is a superoperator which we will denote by dissipator, which describes the incoherent dynamics. In this work, we assume time-local quantum master equation. Such kind of generators can be found as limiting case of a Nakajima-Zwanzig master equation when the environment’s correlation time is much shorter than the system’s characteristic time scale [16]. Being an approximation, it is not warranted that the corresponding map is CPTP. The Lindblad form, as in the model considered in Ref. [5] warrants positivity, but does not deliver the correct asymptotic limit. Other known forms that fulfill the fluctuation-dissipation theorem do not generally garantee positivity. In what follows we investigate how the physical behavior of the quantum active particle depend on these different choices and whether, in the classical limit, they reproduce the classical active dynamics.
@@ II.2 Models for the dissipator
Here, we first list the different kinds of dissipators considered in this work, highlighting their defining features and the motivation behind their choice. Each dissipator provides an approximate description of different types of noise that the system might encounter, reflecting a range of experimental situations and setups. In what follows, we use the notation $\hat{\rho}(x_{c}(t),t)\equiv\hat{\rho}(t)$ for a single stochastic realization $x_{c}(t)$.
II.2.1 Static Lindblad dissipator
We start from the static Lindblad dissipator $\hat{\mathcal{D}}_{L}$ [31],
$\hat{\mathcal{D}}_{L}\equiv\hat{\mathcal{D}}(\hat{\rho}(t))$ $=$ $\frac{\gamma}{2}\bar{n}\left(\hat{a}^{\dagger}\hat{\rho}(t)\hat{a}-\frac{1}{2}\left\{\hat{a}\hat{a}^{\dagger},\hat{\rho}(t)\right\}\right)$
where $\left\{\cdots,\cdots\right\}$ is an anti-commutator defined by $\left\{\hat{A},\hat{B}\right\}=\hat{A}\hat{B}+\hat{B}\hat{A}$. The annihilation $\hat{a}$ and creation operators $\hat{a}^{\dagger}$ are defined as
@@ II.2.1 Static Lindblad dissipator
In the dissipator (II.2.1), $\gamma$ sets the dissipation strength, and $\bar{n}$ is the expectation value of the number operator for a heat bath of temperature $T$,
@@
Such a dissipator has been previously considered in the case of a quantum active particle in [5] for weak dissipation $\gamma$. In what follows, we control the dissipation strength via the gain and loss rates $\nu_{\pm}$, defined as

### v2

@@ Abstract
Active matter denotes a system of particles immersed in an external environment, from which the particles extract energy continuously in order to perform directed motion. Extending the paradigm of active matter to a quantum framework requires an appropriate description of the environment. In this work, we consider a driven quantum particle undergoing noise and dissipation, with external driving exhibiting characteristics of classical activity. We model the non-unitary dynamics with time-local master equations and analyze the particle motion at different time scales for different forms of the master equations, satisfying different criteria. We systematically compare predictions on the dynamics of particle trajectories and thereby we uncover how the particle motion evolves under the interplay of quantum effects, dissipation, and active-like dynamics. These results are essential for guiding possible experiments aimed at realizing quantum analogues of classical active systems.
@@ I Introduction
Quantum active matter is an emerging and rapidly developing field [1, 2, 3, 4, 5, 6, 7] that aims at identifying quantum mechanical dynamics which can be classified as active [8]. In classical physics, activity arises from the local absorption of energy from internal or external sources and its subsequent conversion into persistent, directed motion [9, 10, 11]. Recent works have shown that the concept of active motion can be transferred to quantum systems [2, 3, 4, 7]. A paradigmatic example is a quantum particle bound to a potential, whose center undergoes stochastic fluctuation [5], see Fig. 1. The resulting quantum motion can then exhibit key characteristics of the motion of “self-propelled” particles [8, 9, 11], i.e., ballistic motion at intermediate times and active diffusion at late times, while the system remains unbiased (isotropic) after averaging over many realizations. While the approach suggested in [5] generally works both in Hamiltonian and in dissipative systems, the latter case shows closer analogies with classical active matter and is therefore of particular importance. This brings to the fore the need to understand how to consistently model dissipation in activated quantum systems, which motivates the present study.
Figure 1: Schematic representation of a possible experiment – an external parabolic potential (yellow) captures a quantum particle (blue wave function). The potential moves along a trajectory $x_{c}(t)$ corresponding to colored noise. Since the quantum particle is constantly driven away from the instantaneous trap center, the system never truly reaches equilibrium, and the quantum particle thereby mimics the properties of active motion incorporated into the driving protocol $x_{c}(t)$.
@@
In classical active matter, dissipation is a crucial factor since active matter dynamics and thermodynamics typically arise from the energy exchange with an external environment [12, 13, 14, 15]. A quantum mechanical analogue of active matter shall also possess these features and necessarily calls for an adequate formalism. However, the interaction of a quantum system with an external environment typically tends to induce loss of quantum coherence. Hence, one key question is to understand the role that a reservoir plays for quantum activity.
@@
The theory of open quantum systems [16] provides a sound description of the dynamics of a quantum system that exchanges energy and establishes correlations with an external reservoir. In the presence of interactions with an external reservoir, the evolution of the quantum state is determined by a completely positive and trace-preserving (CPTP) time evolution of the system’s density matrix. The latter describes the dynamics of the quantum probability amplitudes, from which interferences originate, and of the classical probabilities due to partial knowledge about the system, which determine the dynamics in the classical limit.
@@
Generators of the CPTP map can be cast in the form of Nakajima-Zwanzig master equations [17, 18, 19], which can be rigorously derived from a Hamiltonian description of coupled system and external environment by eliminating the environment’s variables from the equations of motion of the system’s variables [16]. They are generally inhomogeneous differential equations for the system’s density operator that are nonlocal in time. Their specific form depends on the initial state of the reservoir, on the initial correlations between system and reservoir, and on the system-reservoir interactions. These equations are tractable in few limits [20] and are usually simplified by means of plausible approximations [21]. One prominent approximation is the Born-Markov one, for which the master equation becomes a homogeneous differential equation that is local in time [16, 22, 21] (see also [23]). Different forms of time-local master equation have been discussed and used in the literature. Some satisfy the CPTP requirement [24, 25, 26]. For instance, Born-Markov master equations derived using the secular approximation [27, 16, 28] and the rotating-wave approximation [29, 28] are well suited for the description of quantum optical systems. Other approaches are designed to be thermodynamically consistent, making them more natural from the perspective of statistical mechanics [30, 31, 32, 33]. For an overview of time-local master equations and an accuracy assessment see [21]. This distinction is not merely theoretical. Different master equations can best describe different regimes and experimental scenarios [34, 22]. Understanding how quantum active dynamics can emerge for different models, describing dissipation and dephasing, can be crucial for experiments aiming at observing quantum active dynamics.
@@
The purpose of this work is to analyze the effect of different types of quantum time-local master equations on the dynamics of quantum active matter. Our starting point is the framework of Ref. [5], where the features of active matter reproduced by a quantum particle confined by a potential and driven with colored noise. In this paper we extend this approach to consider different generators of the open-quantum system dynamics. By systematically comparing several dissipative master equations, we uncover the interplay between quantum effects and active-like dynamics across weak and strong dissipation regimes. Depending on the underlying dissipator, describing the non-unitary part of the dynamics, one may obtain a quantum-active evolution that ensures positivity of the density matrix, or one that reproduces the standard active matter dynamics in the classical limit – but not both simultaneously.
The paper is organized as follows: in Sec. II we introduce the models and discuss the details of the numerical procedure. In Sec. III we provide the results for the different master equations and discuss them. In Sec. IV we draw our conclusions. Details of theoretical models and analysis are provided in the Appendices.
II Open quantum system description of a quantum active particle
The purpose of this section is to introduce models of a quantum active particle in the presence of an external environment.
@@ II.1 Model
Our starting point is the minimal model of quantum active matter of Ref. [5], which is inspired by the active Ornstein-Uhlenbeck particle (AOUP) and is one dimensional. In the classical AOUP model, the particle’s self-propulsion arises from colored noise $x_{c}$, which drives persistent active motion [35, 36, 37, 38, 39, 40, 41]. Specifically, the variable $x_{c}(t)$ is the classical trajectory of a particle that solves the equation of motion
@@
where $\tau$ is the inverse damping rate (or equivalently, the persistence time), $D$ is the diffusion coefficient in units of the damping rate, and $\eta(t)$ is a Gaussian white noise with zero mean and correlations $\overline{\eta(t^{\prime})\eta(t^{\prime\prime})}=\delta(t^{\prime}-t^{\prime\prime})$.
In the quantum-mechanical model proposed in Ref. [5], a quantum particle of mass $m$ is confined by a time-dependent harmonic potential. The potential has a fixed frequency $\omega$ but the center follows the prescribed Ornstein-Uhlenbeck trajectory $x_{c}(t)$. Thereby, the quantum particle is governed by a time-dependent Hamiltonian of the form
@@
where $x_{c}(t)$ is a classical variable undergoing the dynamics of Eq. (1); $\hat{x}$ and $\hat{p}$ are the canonically-conjugated position and momentum along $x$, satisfying the commutation relations $\hat{x},\hat{p}]=i\hbar$, where $\hbar$ is Planck’s constant. The classical particle establishes a contact between the quantum particle and a thermal bath. Since the external drive $x_{c}(t)$ is governed by the classical equation of motion (1) and unaffected by the state of the quantum particle, we can represent the state of the quantum system by a probability distribution over quantum trajectories $|\psi(x_{c}(t),t)\rangle$. Here, $|\psi(x_{c}(t),t)\rangle$ is the solution of the Schrödinger equation $\partial_{t}|\psi(x_{c}(t),t)\rangle=\hat{H}(x_{c}(t))|\psi(x_{c}(t),t)\rangle$ for a given trajectory $x_{c}(t)$. The state of the quantum particle is given by the density matrix $\hat{\varrho}(t)$
$\hat{\varrho}(t)=\int Dx_{c}(t)\mathcal{P}[x_{c}(t)]|\psi(x_{c}(t),t)\rangle\langle\psi(x_{c}(t),t)|\,$, (3)
where $P[x_{c}]$ denotes the probability functional that the particle follows the trajectory $x_{c}(t)$ and the functional integral is performed over all paths, such that $\int Dx_{c}(t)\mathcal{P}[x_{c}(t)]=1$.
In this work we assume that the quantum active particle is additionally coupled to a reservoir. The motion results from the interplay of the energy exchange of the quantum particle with an external reservoir and the coupling with the classical active particle, undergoing the dynamics governed by Eq. (1). For a given trajectory $x_{c}(t)$, the state of the system is now described by a density operator $\hat{\rho}(x_{c}(t),t)$. Since $x_{c}(t)$ represents the stochastic trajectory of a classical degree of freedom, physical observables must be averaged not only over quantum states but also over the ensemble of possible realizations of $x_{c}(t)$. In other words, the quantum system is subject to a noisy classical driving, and its effective behavior emerges only after performing both the quantum and the classical ensemble average. In general, the state $\hat{\rho}(x_{c}(t),t)$ is evolved from the initial state $\hat{\rho}(x_{c}(0),0)$ by the CPTP map $\hat{\rho}(x_{c}(t),t)=\hat{\Lambda}_{x_{c}(t)}[\hat{\rho}(x_{c}(0),0)]$. By averaging over all trajectories, the state of the system is now given by
@@
In order to determine the dissipative dynamics, we consider generators that are local in time, such that $\partial_{t}\hat{\rho}={\mathcal{L}}(x_{c}(t),t)\hat{\rho}$. The superoperator $\mathcal{L}$} has the form
${\mathcal{L}}(x_{c}(t),t)\hat{\rho}$
$=$ $-\frac{i}{\hbar}\left[\hat{H}(x_{c}(t)),\hat{\rho}(x_{c}(t),t)\right]+{\mathcal{D}}(x_{c}(t),\hat{\rho}(x_{c}(t),t)),$
where $\left[\cdot,\cdot\right]$ is the commutator $\left[\hat{A},\hat{B}\right]=\hat{A}\hat{B}-\hat{B}\hat{A}$ and $\mathcal{D}$} describes the incoherent dynamics, which in what follows we denote by “dissipator”. In this work, we assume time-local quantum master equation, which can be found as limiting case of a Nakajima-Zwanzig master equation when the environment’s correlation time is much shorter than the system’s characteristic time scale [16]. Being an approximation, it is not warranted that the corresponding map is CPTP. Positivity is warranted by the Lindblad form, as the one in the model of Ref. [5]. However, the Lindblad master equation does not deliver the correct asymptotic limit [29, 42]. Other known forms satisfy the fluctuation-dissipation theorem but do not generally guarantee positivity [43, 44]. In what follows, we investigate how the physical behavior of the quantum active particle depends on these different choices.
@@ II.2 Models for the dissipator
Here, we first introduce the different kinds of dissipators considered in this work, highlighting their defining features and the motivation behind their choice. Each dissipator provides an approximate description of different types of noise and of the regimes, possibly reflecting a range of experimental situations and setups. In what follows, we use the notation $\hat{\rho}(x_{c}(t),t)\equiv\hat{\rho}(t)$ for a single stochastic realization $x_{c}(t)$.
II.2.1 Lindblad dissipator
We consider a form of Lindblad dissipator, inspired by the one used in the model of a quantum active particle of Ref. [5]. The idea of [5] was to mimic active motion of a quantum particle by means of a trapping potential whose center follows the noisy trajectory of a classical particle. The quantum particle is also exposed to a thermal bath, whose effect is a weak perturbation with respect to the noise mediated by the classical particle. One issue of the form of Lindblad dissipator of [5] is that it does not preserve physical properties at large thermalization rates, see the discussion in Appendix A.2. In this work we consider a modification, that describes a thermal bath contact moving with $x_{c}(t)$. Before we give its explicit form, we introduce the annihilation and creation operators of the harmonic oscillator, $\hat{\tilde{a}}(t)$ and $\hat{\tilde{a}}^{\dagger}(t)$, such that Hamiltonian (2) takes the form $H=\hbar\omega(\hat{\tilde{a}}^{\dagger}(t)\hat{\tilde{a}}(t)+1/2$). These operators are related to the position and momentum operators of the oscillator centered at the origin, $\hat{x}=x_{*}(\hat{a}+\hat{a}^{\dagger})$ and $\hat{p}=-i(\hat{a}-\hat{a}^{\dagger})\hbar/(2x_{*})$ with $x_{*}=\sqrt{\hbar/(2m\omega)}$, by the relations
@@ II.2.1 Lindblad dissipator
The Lindblad dissipator takes the form (see Appendix B for details of the derivation):
${\mathcal{D}}_{L}$ $=$ $\frac{\gamma}{2}\bar{n}\left(\hat{\tilde{a}}^{\dagger}(t)\hat{\rho}(t)\hat{\tilde{a}}(t)-\frac{1}{2}\left\{\hat{\tilde{a}}(t)\hat{\tilde{a}}^{\dagger}(t),\hat{\rho}(t)\right\}\right)$
$+$ $\frac{\gamma}{2}(\bar{n}+1)\left(\hat{\tilde{a}}(t)\hat{\rho}(t)\hat{\tilde{a}}^{\dagger}(t)-\frac{1}{2}\left\{\hat{\tilde{a}}^{\dagger}(t)\hat{\tilde{a}}(t),\hat{\rho}(t)\right\}\right),$
where $\left\{\hat{A},\hat{B}\right\}=\hat{A}\hat{B}+\hat{B}\hat{A}$ for any two operators $\hat{A}$ and $\hat{B}$ acting in the Hilbert space of the quantum particle. We note that for the mathematical consistency of the suggested dissipator, $\dot{x}_{c}(t)$ must be a continuous function so that the driving protocol is continuous and differentiable in time (the necessity of this condition is discussed in Appendix C). The process $x_{c}(t)$ set by Eq. (1) satisfies this condition.
Dissipator (II.2.1) describes the dynamics emerging from the thermal contact of an oscillator with a thermal bath at temperature $T$. Here, $\bar{n}$ is the expectation value of the oscillator number operator at thermal equilibrium,
@@
where $k_{B}$ is the Boltzmann constant, and $\gamma$ is the thermalization rate [45].
In this work we will analyze the interplay between the classical colored noise induced by the fluctuations of the potential and thermalization for thermalization rates. In what follows, we rewrite the dissipation strength using the gain and loss rates $\nu_{\pm}$, defined as